I need help to write an equation of a line perpendicular to the line 12y+15x=6
Answer by nabla(475)

Solve for y.
12y+15x=6
y=(-15x+6)/12
The slope is obviously -15/12.
A line perpendicular to that line will have a slope that is the negative reciprocal of that slope.
(-1)/(-15/12)=12/15.
A line would be one such as y=(12/15)x+1